# Breast Cancer Diagnosis: A Comprehensive Exploration of Explainable Artificial Intelligence (XAI) Techniques

### v1

@@ 3 Dataset and Modalities for Breast Cancer
Features {BreakHis [46] and others}, Mammograms {Breast Cancer Dataset (BCD) [47], Mammographic Mass (MM) [48], Digital Database for Screening Mammography (DDSM) [49], INbreast [50], Wisconsin Breast Cancer (WBC) [51], and Radiological Society of North America (RSNA) [52] and others}, Genes {The Cancer Genome Atlas (TCGA) and Gene Expression Omnibus (GEO) (TCG-GEO) [53, 54], and National Center for Biotechnology Information (NCBI) and Gene Expression Omnibus (NCBI-GEO) [55, 56]}, Ultrasound {Breast Ultrasound Image (PlumX) [57], Breast Ultrasound Image (BUSI) [58], and others}, Whole Slide Images (WSI) {ITU-MED-1, ITU-MED-2 [22], and others}, Digital Breast Tomosynthesis (DBT) {Two in House (TiH) [24] and others}, Computed Tomography (CT){others}, and Infrared {Mastology DRM (MDRM) [59]}. Others: Proprietary datasets.
@@
The literature review has indicated that XAI methods have been significantly employed across various modalities, with the features dataset emerging as a prominently utilized resource for enhancing the transparency and interpretability of models in breast cancer research, as exhibited in Figure 5. A feature dataset in breast cancer research encompasses specific attributes gathered from various diagnostic tools, including imaging, genetic tests, and biochemical analysis. Imaging features from mammograms detail tumor characteristics like shape and texture [60]. Histopathological features involve the microscopic analysis of tissue samples [29], and genomic features highlight gene expression and mutations [21],  [35][16],  [44]. Biochemical features such as levels of enzymes, including lactate dehydrogenase, alkaline phosphatase, and aromatase, as well as hormone assays, provide insights into biological activities associated with cancer [28],  [38]. Clinical features integrate personal data such as age and family history[40][36].
@@ 5.2 Class Activation Map and Gradient-weighted Class Activation Map (Grad-CAM)
Where $L_{Grad-CAM}^{c}$ represents the Grad-CAM localization map for class $c$, $A^{k}$ is the $k-th$ feature map of the convolutional layer and $\alpha_{k}^{c}$ is the importance weight of the $k$-th feature map for class $c$.
@@ 6 Discussion
Shapley values also offer three useful properties that are not found in other XAI methods [65][129]. First, SHAP values provide local accuracy, meaning they accurately capture the difference between the expected output of the model and the output for a specific instance of input data. This means they provide insights into how individual features contribute to each prediction. Second, if a feature is missing (i.e., not present in the input data), its attribution value in the SHAP analysis is set to zero, ensuring that the absence of a feature is appropriately accounted for in the explanation. Third, changes in the model should not result in contradictory changes to the SHAP values assigned to individual features. If increasing the importance of a feature in the model’s prediction also increases its corresponding SHAP value, then consistency is maintained. These advantageous properties make the SHAP model an ideal choice for different research studies’ objectives. It offers a robust and reliable method for explaining model predictions and comparing different models, regardless of their type or complexity. For example, SHAP performed well with the XGBoost model, as witnessed in  [28][38][40][36], and [43].
@@
Grad-CAM is a widely applied XAI method in medical images [131]. It generates heatmaps that highlight regions of an image that contribute significantly to the model’s decision, providing insights into which features or areas the model focuses on for classification. This interpretability aids radiologists or clinicians in understanding and validating the model’s predictions. The research studies discussed in this survey applied Grad-CAM to predict whether an image contains cancerous features and localize the specific areas in the image that contribute to the prediction. This localization capability can help pinpoint the location of abnormalities within breast tissue, potentially improving diagnostic accuracy. Techniques such as [13][14][20], and [25] indicated that Grad-CAM serves as a valuable tool to corroborate the findings of deep learning models by providing visual evidence of the features that drive the predictions.

### v2

@@ 3 Dataset and Modalities for Breast Cancer
Features {BreakHis [46] and others}, Mammograms {Breast Cancer Dataset (BCD) [47], Mammographic Mass ($M$) [48], Digital Database for Screening Mammography (DDSM) [49], INbreast [50], Wisconsin Breast Cancer (WBC) [51], and Radiological Society of North America (RSNA) [52] and others}, Genes {The Cancer Genome Atlas (TCGA) and Gene Expression Omnibus (GEO) (TCG-GEO) [53, 54], and National Center for Biotechnology Information (NCBI) and Gene Expression Omnibus (NCBI-GEO) [55, 56]}, Ultrasound {Breast Ultrasound Image (PlumX) [57], Breast Ultrasound Image (BUSI) [58], and others}, Whole Slide Images (WSI) {ITU-MED-1, ITU-MED-2 [22], and others}, Digital Breast Tomosynthesis (DBT) {Two in House (TiH) [24] and others}, Computed Tomography (CT){others}, and Infrared {Mastology DRM (MDRM) [59]}. Others: Proprietary datasets.
@@
The literature review has indicated that XAI methods have been significantly employed across various modalities, with the features dataset emerging as a prominently utilized resource for enhancing the transparency and interpretability of models in breast cancer research, as exhibited in Figure 5. A feature dataset in breast cancer research encompasses specific attributes gathered from various diagnostic tools, including imaging, genetic tests, and biochemical analysis. Imaging features from mammograms detail tumor characteristics like shape and texture [60]. Histopathological features involve the microscopic analysis of tissue samples [29], and genomic features highlight gene expression and mutations [21], [35], [16], [44]. Biochemical features such as levels of enzymes, including lactate dehydrogenase, alkaline phosphatase, and aromatase, as well as hormone assays, provide insights into biological activities associated with cancer [28], [38]. Clinical features integrate personal data such as age and family history[40], [36].
@@ 5.2 Class Activation Map and Gradient-weighted Class Activation Map (Grad-CAM)
Where $L_{Grad-CAM}^{c}$ represents the Grad-CAM localization map for class $c$, $A^{k}$ is the kthk-th feature map of the convolutional layer and $\alpha_{k}^{c}$ is the importance weight of the $k$-th feature map for class $c$.
@@ 6 Discussion
Shapley values also offer three useful properties that are not found in other XAI methods [65], [129]. First, SHAP values provide local accuracy, meaning they accurately capture the difference between the expected output of the model and the output for a specific instance of input data. This means they provide insights into how individual features contribute to each prediction. Second, if a feature is missing (i.e., not present in the input data), its attribution value in the SHAP analysis is set to zero, ensuring that the absence of a feature is appropriately accounted for in the explanation. Third, changes in the model should not result in contradictory changes to the SHAP values assigned to individual features. If increasing the importance of a feature in the model’s prediction also increases its corresponding SHAP value, then consistency is maintained. These advantageous properties make the SHAP model an ideal choice for different research studies’ objectives. It offers a robust and reliable method for explaining model predictions and comparing different models, regardless of their type or complexity. For example, SHAP performed well with the XGBoost model, as witnessed in [28], [38], [40], [36], and [43].
@@
Grad-CAM is a widely applied XAI method in medical images [131]. It generates heatmaps that highlight regions of an image that contribute significantly to the model’s decision, providing insights into which features or areas the model focuses on for classification. This interpretability aids radiologists or clinicians in understanding and validating the model’s predictions. The research studies discussed in this survey applied Grad-CAM to predict whether an image contains cancerous features and localize the specific areas in the image that contribute to the prediction. This localization capability can help pinpoint the location of abnormalities within breast tissue, potentially improving diagnostic accuracy. Techniques such as [13], [14], [20], and [25] indicated that Grad-CAM serves as a valuable tool to corroborate the findings of deep learning models by providing visual evidence of the features that drive the predictions.